OpenStudy (anonymous):

Medal Reward! . Which shows the prime factors of 12? 2 × 2 × 5 2 x 2 x 3 2 x 3 x 5 2 x 6

OpenStudy (cggurumanjunath):

2*2*3

OpenStudy (cggurumanjunath):

2*2*5=4*5=20 2*2*3=4*3=12 2*3*5=6*5=30 2*6=12

OpenStudy (cggurumanjunath):

2*6=2*2*3

OpenStudy (anonymous):

12 / \ 6 2 / \ 3 2 2*3*2

OpenStudy (cggurumanjunath):

2*2*3=12 IS THE RIGHT ANSWER.